# On the numerically and cohomologically trivial automorphisms of elliptic surfaces II: $\chi(S)>0$

### v1

@@ 1.2. Organization of the paper
Section 7 introduces our main technique to construct explicit properly elliptic surfaces with given groups of numerically trivial automorphism.
@@
Convention 1.6.
Throughout the paper, we work over the complex numbers, although our constructions work over arbitrary fields as long as the characteristic does not divide the group order.
@@ 2.1. Special subgroups of the automorphism groups
For a properly elliptic surface $S$, the existence of vector fields is equivalent to being pseudo-elliptic, which means that $S=(C\times E)/G$ is isogenous to a product, and with $G$ acting by translations on the elliptic curve $E$: in particular $\chi(S)=0$, and we have the first of the two cases studied in Part I, [CFGLS24].
@@ 2.5. Local behavior of an automorphism around an $m\mathrm{I}_{k}$ fibre
If in fact $\sigma$ acts on the i-th component via $\beta_{i}$, then, since $\sigma(z)=\tau(z)=\epsilon^{m}z$, in view of the fact that in local coordinates at a node of $(F_{0})_{\mathrm{red}}$ it acts via multiplication by $(\beta_{i+1},\beta_{i}^{-1})$ we reach a contradiction by assuming that $\beta_{i+1}=\beta_{i}=1$.
@@ 3.3. Cohomologically trivial automorphisms
$\Lambda^{\prime}=U^{\prime}\oplus A_{2}\oplus E_{6}\subset\Lambda=Num(S)=H^{2}(S,\mathbb{Z})/\operatorname{Tors}$
@@
Here $C_{2}$ has multiplicity two as fibre component and the self-intersection numbers are indicated to the left. Contracting first $C_{0}$ and then the image of $C_{2}$, we obtain the desired elliptic surface $S$ with a fibre of type $\mathrm{III}$ (and $S$ is Enriques, if $s=1$, and properly elliptic for $s>1$). This admits two natural bisections $O_{S},R_{S}$ induced by the 2-torsion of $E$. On the singular model $S_{0}$, $R_{S}$ meets exactly the singularities of type $A_{1}$ (induced by the 2-torsion points $\frac{1}{2},\frac{i}{2}$), and $O_{S}$ the other two, induced by 00 and $t_{0}$. On $S$, it follows that $O_{S}$ meets the $\mathrm{III}^{*}$ fiber transversally in the simple components $\Theta_{0},\Theta_{6}$, and $\mathrm{III}$ in one point (different from the node) on each component $C_{1},C_{3}$, while $R_{S}$ intersects $\mathrm{III}$ in the node (which $C_{2}$ was contracted to) and $\mathrm{III}^{*}$ in the central double component $\Theta_{7}$, as depicted in the following dual graph:
@@ 4. Jacobian elliptic fibrations and their Mordell–Weil groups
As mentioned in the introduction, the function field extension $\mathbb{C}(S)\supset\mathbb{C}(B)$ provides a curve ${\mathcal{C}}$ of genus $1$ (the fibre of $f$ over the generic point of $B$) defined over the non algebraically closed field ${\mathcal{K}}:=\mathbb{C}(B)$.
And the subgroup of $Aut(S)$ of automorphisms preserving the fibration and inducing the identity on $B$ is isomorphic to the subgroup $\mathrm{Aut}_{{\mathcal{K}}}({\mathcal{C}})$ of the automorphisms of ${\mathcal{C}}$ which act as the identity on the base field ${\mathcal{K}}$.
@@
We denote as usual $\mathrm{Jac}({\mathcal{C}}):=\operatorname{Pic}^{0}({\mathcal{C}})$, and ${\mathcal{C}}$ is a principal homogeneous space over $\mathrm{Jac}({\mathcal{C}})$ ([Shaf-65], Chapter 2 of [Dolg81]), in particular, the Mordell–Weil group, the group of ${\mathcal{K}}$-rational points of $\mathrm{Jac}({\mathcal{C}})$ acts on ${\mathcal{C}}$.
@@
For the other inclusion, take any $\sigma\in\mathrm{Aut}_{B,{\mathcal{O}}}(S)$. Assuming that $\sigma$ is not a translation on a general fibre $F$, it necessarily fixes some points on $F$ and is thus of finite order. Hence we can argue with the quotient surface $S/\langle\sigma\rangle$, which is a $\mathbb{P}^{1}$-fibration over $B$ whence $p_{g}(S/\langle\sigma\rangle)$ vanishes. On the hand, by the choice of $\sigma$, we have $p_{g}(S/\langle\sigma\rangle)=p_{g}(S)>0$. This gives the required contradiction.
@@
It follows that $\mathrm{Aut}_{B,\mathbb{Q}}(X)$ fixes (pointwise) each section of $h$. Since $\mathrm{Aut}_{B,\mathbb{Q}}(X)$ is nontrivial group, the Mordell–Weil group $\mathrm{MW}(X)$ of sections is a finite group.
(2) If $p_{g}(X)>0$, then $\mathrm{Aut}_{\mathbb{Q}}(X)\subseteq t(\mathrm{MW}(X)_{\mathrm{tor}})=t(\mathrm{MW}(X))$ by Lemma 4.4 and (1). But then any torsion section $P$ meets some fibre in a non-identity component; this is a consequence of (2.10). It follows that $t(P)$ acts as a non-trivially permutation on the (simple) fibre components, hence $t(P)\not\in\mathrm{Aut}_{\mathbb{Q}}(X)$.
@@ 5. Bounding the torsion part of the Mordell–Weil groups
Let $B(N)$ and $B_{1}(N)$ be the modular curves associated to $\Gamma(N)$ and $\Gamma_{1}(N)$ respectively, and let $g(N)$ and $g_{1}(N)$ be their respective the genera. For $N\geq 5$, we have, by [Shi72, (5.3)] and [Mar05, page 314, line 2],
@@ 7. The main construction
Let $h\colon X\rightarrow B$ be a jacobian elliptic fibration with zero section $O$, and $G\subset\mathrm{MW}(X)_{\mathrm{tor}}$ a subgroup. If $G\cong\mathbb{Z}/N\mathbb{Z}\times\mathbb{Z}/N^{\prime}\mathbb{Z}$ such that at every prime $p$, the $p$-adic valuations satisfy $v_{p}(N)=v_{p}(N^{\prime})$ or $\max(v_{p}(N),v_{p}(N^{\prime}))\leq 1$, then $G$ splits at every singular fibre of type $\mathrm{I}_{n}$.
@@
$\iota(S_{p}(G))\cong\mu_{p^{v_{p}}}\times\mathbb{Z}/{p^{v_{p}}}\mathbb{Z}$
for size reasons. Meanwhile, if $v_{p}(N)=1$ and $v_{p}(N^{\prime})=0$, then, using the notation of Definition 7.1, either $\iota(S_{p}(G))\cap\mathbb{C}^{*}$ or $\mathrm{im}(\iota(S_{p}(G))\rightarrow\mathbb{Z}/n\mathbb{Z})$ is trivial; in other words, we have
(7.2) $\iota(S_{p}(G))=\{1\}\times\mathbb{Z}/p\mathbb{Z},\quad\text{ or }\quad\iota(S_{p}(G))=\mu_{p}\times\{0\}.$
where the second alternative may only persist if $p\mid n$. For symmetry reasons, this completes the proof. ∎
@@
The bounds follow along the same lines as in the proof of Proposition 5.5. Namely, with $G\subset\mathrm{MW}(J(S))$ of size $|G|>4$, $J(S)$ is a base change of $X(G)$, and since $S$ and $J(S)$ share the same invariants as discussed in 2.4, we infer that
@@
10. Cohomologically trivial automorphisms – Proof of 2-part of Theorem 1.4
@@
11. Cohomologically trivial automorphisms – Proof of 3-part of Theorem 1.4
@@
from 2.4 where the trivial lattice Triv$(X)$ is generated by fibre components and zero section (in parallel to the argument around (11.2)), or from considering explicit equations as in [Bea82], [MP86] or [SS19, Table 8.3].
@@
which parallels the trivial lattice of a jacobian elliptic surface reviewed in 2.4. Note that integral overlattices of $L$ are in bijective correspondence with isotropic subgroups of the discriminant group $L^{\vee}/L$. Presently, this group is isomorphic to $A_{8}^{\vee}/A_{8}=\langle\Theta_{1}^{\vee}\rangle\cong\mathbb{Z}/9\mathbb{Z}$, with generator the dual vector $\Theta_{1}^{\vee}$ of square $-8/9$. Since $\mathrm{Num}(S)$ is unimodular, we infer that it is obtained from $L$ by adjoining the class of the dual vector

### v2

@@ 1.4. Main results for cohomologically trivial automorphisms
A posteriori, we can infer the following from our results in this paper combined with those from [CFGLS24]:
Theorem 1.6.
For any algebraic properly elliptic surface $S$, $\mathrm{Aut}_{\mathbb{Z}}(S)$ is abelian.
@@ 1.5. Organization of the paper
Section 7 introduces our main technique to construct explicit properly elliptic surfaces with given groups of numerically trivial automorphisms.
@@
Convention 1.7.
Throughout the paper, we work with algebraic surfaces over the complex numbers, although our constructions work over arbitrary fields as long as the characteristic does not divide the group order.
@@ 2.1. Special subgroups of the automorphism groups
For an algebraic properly elliptic surface $S$, the existence of vector fields is equivalent to being pseudo-elliptic, which means that $S=(C\times E)/G$ is isogenous to a product, and with $G$ acting by translations on the elliptic curve $E$: in particular $\chi(S)=0$, and we have the first of the two cases studied in Part I, [CFGLS24].
@@ 2.2. Notation and basic formulae
2.3. Properly elliptic surfaces with $\chi(S)>0$ are Kähler, but not necessarily algebraic
Here we borrow first from the survey paper [Cat85], especially the classification table on page 21, remarks 44, 45 on page 22, and the discussion on pages 25-28, giving a very succinct presentation of the results obtained by Kodaira in [Kod60], [Kod63], [Kod64].
If a compact complex surface $S$ has $Kod(S)=1$, then we are in one of the two classes $4_{1}$ or $6_{1}$.
The latter class $6_{1}$ consists of elliptic quasi-bundles, and for these $b_{1}(S)$ is odd, hence they are not Kähler, and obviously for them $\chi(S)=0$.
In the class $4_{1}$, $b_{1}(S)$ is even, the algebraic dimension $a(S)\in\{1,2\}$, $S$ deforms to an algebraic surface and is Kähler.
Hence
I) a surface with $Kod(S)=1$ and $\chi(S)>0$ is necessarily Kähler.
Moreover
II) If $p_{g}(S)=0$ and $b_{1}(S)$ is even, then $S$ is algebraic (since $b^{+}(S)=2p_{g}(S)+1$).
If $S$ is not algebraic, then it has algebraic dimension $a(S)=1$ and the pluricanonical elliptic fibration $f:S\rightarrow C$ does not admit any multisection.
These surfaces $S$ are obtained from a fibration without multiple fibres as quotient by an abelian group $G$ operating equivariantly on $f^{\prime}:S^{\prime}\rightarrow B$: meaning that $S=S^{\prime}/G$, $C=B/G$.
In particular we have that $\chi(S^{\prime})>0\Leftrightarrow\chi(S)>0$.
If $f:S\rightarrow C$ has no multiple fibres and all fibres are smooth, then $S$ is a deformation of a product $E\times C$, where $E$ is an elliptic curve.
Assume now that $p_{g}(S)>0$ and $a(S)=1$.
As proven by Kodaira in [Kod60], $S^{\prime}$ is a deformation of an elliptic surface $J\rightarrow B$ admitting a section, which is called the Jacobian elliptic surface of $S^{\prime}$.
For a fixed Jacobian elliptic surface $J$, the parameter space for the surfaces $S^{\prime}$ is the cohomology group $H^{1}(C,{\mathcal{B}})$ , ${\mathcal{B}}$ being the sheaf of local holomorphic sections of the base curve $B$ to the open set of $J$ where the fibre is smooth. And to $S^{\prime}$ corresponds the invariant $\eta\in H^{1}(C,{\mathcal{B}})$.
We borrow now from [FM94] Proposition 5.23:
III) $S^{\prime}$, hence $S$, is algebraic if and only if $\eta$ is a torsion element.
IV) If $S^{\prime}$, equivalently $J$, has some singular fibres, then
$H^{1}(B,{\mathcal{B}})=H^{2}(\mathcal{O}_{J})/H^{2}(J,{\mathbb{Z}}).$
The above criterion shows that, if $p_{g}(S)>0$, and there are no multiple fibres, then we get a dense set of algebraic surfaces which deform to $S$. And also that there exist surfaces which are not algebraic and which have $Kod(S)=1$ and $\chi(S)>0$.
We shall consider in the sequel to this article the non algebraic case, the main difficulty being that the Jacobian of an elliptic surface with multiple fibres has been defined only for algebraic elliptic surfaces.
@@ 2.6. Local behavior of an automorphism around an $m\mathrm{I}_{k}$ fibre
If in fact $\sigma$ acts on the $i$th component via $\beta_{i}$, then, since $\sigma(z)=\tau(z)=\epsilon^{m}z$, in view of the fact that in local coordinates at a node of $(F_{0})_{\mathrm{red}}$ it acts via multiplication by $(\beta_{i+1},\beta_{i}^{-1})$ we reach a contradiction by assuming that $\beta_{i+1}=\beta_{i}=1$.
@@ 3.3. Cohomologically trivial automorphisms
$\Lambda^{\prime}=U^{\prime}\oplus A_{2}\oplus E_{6}\subset\Lambda=\mathrm{Num}(S)=H^{2}(S,\mathbb{Z})/\operatorname{Tors}$
@@
Here $C_{2}$ has multiplicity two as fibre component and the self-intersection numbers are indicated to the left. Contracting first $C_{0}$ and then the image of $C_{2}$, we obtain the desired elliptic surface $S$ with a fibre of type $\mathrm{III}$ (and $S$ is Enriques, if $s=1$, and properly elliptic for $s>1$). This admits two natural bisections $O_{S},R_{S}$ induced by the 2-torsion of $E$. On the singular model $S_{0}$, $R_{S}$ meets exactly the singularities of type $A_{1}$ (induced by the 2-torsion points $\frac{1}{2},\frac{i}{2}$), and $O_{S}$ the other two, induced by 00 and $t_{0}$. On $S$, it follows that $O_{S}$ meets the $\mathrm{III}^{*}$ fibre transversally in the simple components $\Theta_{0},\Theta_{6}$, and $\mathrm{III}$ in one point (different from the node) on each component $C_{1},C_{3}$, while $R_{S}$ intersects $\mathrm{III}$ in the node (which $C_{2}$ was contracted to) and $\mathrm{III}^{*}$ in the central double component $\Theta_{7}$, as depicted in the following dual graph:
@@ 4. Jacobian elliptic fibrations and their Mordell–Weil groups
As mentioned in the introduction, the function field extension $\mathbb{C}(S)\supset\mathbb{C}(B)$ provides a curve ${\mathcal{C}}$ of genus $1$ (the fibre of $f$ over the generic point of $B$) defined over the non algebraically closed field ${\mathcal{K}}:=\mathbb{C}(B)$. The subgroup of $\mathrm{Aut}(S)$ of automorphisms preserving the fibration and inducing the identity on $B$ is isomorphic to the subgroup $\mathrm{Aut}_{{\mathcal{K}}}({\mathcal{C}})$ of the automorphisms of ${\mathcal{C}}$ which act as the identity on the base field ${\mathcal{K}}$.
@@
We denote as usual $\mathrm{Jac}({\mathcal{C}}):=\operatorname{Pic}^{0}({\mathcal{C}})$, and ${\mathcal{C}}$ is a principal homogeneous space over $\mathrm{Jac}({\mathcal{C}})$ ([Shaf65], Chapter 2 of [Dolg81]), in particular, the Mordell–Weil group, the group of ${\mathcal{K}}$-rational points of $\mathrm{Jac}({\mathcal{C}})$ acts on ${\mathcal{C}}$.
@@
For the other inclusion, take any $\sigma\in\mathrm{Aut}_{B,{\mathcal{O}}}(S)$. Assuming that $\sigma$ is not a translation on a general fibre $F$, it necessarily fixes some points on $F$ and is thus of finite order. Hence we can argue with the quotient surface $S/\langle\sigma\rangle$, which is a $\mathbb{P}^{1}$-fibration over $B$ whence $p_{g}(S/\langle\sigma\rangle)$ vanishes. On the other hand, by the choice of $\sigma$, we have $p_{g}(S/\langle\sigma\rangle)=p_{g}(S)>0$. This gives the required contradiction.
@@
It follows that $\mathrm{Aut}_{B,\mathbb{Q}}(X)$ fixes (pointwise) each section of $h$. Since $\mathrm{Aut}_{B,\mathbb{Q}}(X)$ is a nontrivial group, the Mordell–Weil group $\mathrm{MW}(X)$ of sections is a finite group.
(2) If $p_{g}(X)>0$, then $\mathrm{Aut}_{\mathbb{Q}}(X)\subseteq t(\mathrm{MW}(X)_{\mathrm{tor}})=t(\mathrm{MW}(X))$ by Lemma 4.4 and (1). But then any torsion section $P$ meets some fibre in a non-identity component; this is a consequence of (2.10). It follows that $t(P)$ acts as a non-trivial permutation on the (simple) fibre components, hence $t(P)\not\in\mathrm{Aut}_{\mathbb{Q}}(X)$.
@@ 5. Bounding the torsion part of the Mordell–Weil groups
Let $B(N)$ and $B_{1}(N)$ be the modular curves associated to $\Gamma(N)$ and $\Gamma_{1}(N)$ respectively, and let $g(N)$ and $g_{1}(N)$ be their respective genera. For $N\geq 5$, we have, by [Shi72, (5.3)] and [Mar05, page 314, line 2],
@@ 7. The main construction
Let $h\colon X\rightarrow B$ be a jacobian elliptic fibration with zero section $O$, and $G\subset\mathrm{MW}(X)_{\mathrm{tor}}$ a subgroup. If $G\cong\mathbb{Z}/N\mathbb{Z}\times\mathbb{Z}/N^{\prime}\mathbb{Z}$ is such that at every prime $p$, the $p$-adic valuations satisfy $|v_{p}(N)-v_{p}(N^{\prime})|\leq 1$, then $G$ splits at every singular fibre of type $\mathrm{I}_{n}$.
@@
(7.2) $\iota(S_{p}(G))\cong\mu_{p^{v_{p}}}\times\mathbb{Z}/{p^{v_{p}}}\mathbb{Z}$
for size reasons. Meanwhile, if $\{v_{p}(N),v_{p}(N^{\prime})\}=\{v_{p},v_{p}+1\}$, then (7.2) becomes a proper inclusion $\iota(S_{p}(G))\supsetneq\mu_{p^{v_{p}}}\times\mathbb{Z}/{p^{v_{p}}}\mathbb{Z}$. Moreover, using the notation of Definition 7.1, either $\iota(S_{p}(G))\cap\mathbb{C}^{*}$ or $\mathrm{im}(\iota(S_{p}(G))\rightarrow\mathbb{Z}/n\mathbb{Z})$ has order $p^{v_{p}}$, and the other group has order $p^{v_{p}+1}$; in other words, we have
(7.3) $\iota(S_{p}(G))\cong\mu_{p^{v_{p}}}\times\mathbb{Z}/p^{v_{p}+1}\mathbb{Z},\quad\text{ or }\quad\iota(S_{p}(G))\cong\mu_{p^{v_{p}+1}}\times\mathbb{Z}/p^{v_{p}}\mathbb{Z},$
where the first alternative may only persist if $v_{p}(n)\geq v_{p}+1$. This completes the proof. ∎
@@
The bounds follow along the same lines as in the proof of Proposition 5.5. Namely, with $G\subset\mathrm{MW}(J(S))$ of size $|G|>4$, $J(S)$ is a base change of $X(G)$, and since $S$ and $J(S)$ share the same invariants as discussed in 2.5, we infer that
@@
10. Cohomologically trivial automorphisms – Proof of the order 2-statement of Theorem 1.4
@@
11. Cohomologically trivial automorphisms – Proof of the order 3-statement of Theorem 1.4
@@
from 2.5 where the trivial lattice Triv$(X)$ is generated by fibre components and zero section (in parallel to the argument around (11.2)), or from considering explicit equations as in [Bea82], [MP86] or [SS19, Table 8.3].
@@
which parallels the trivial lattice of a jacobian elliptic surface reviewed in 2.5. Note that integral overlattices of $L$ are in bijective correspondence with isotropic subgroups of the discriminant group $L^{\vee}/L$. Presently, this group is isomorphic to $A_{8}^{\vee}/A_{8}=\langle\Theta_{1}^{\vee}\rangle\cong\mathbb{Z}/9\mathbb{Z}$, with generator the dual vector $\Theta_{1}^{\vee}$ of square $-8/9$. Since $\mathrm{Num}(S)$ is unimodular, we infer that it is obtained from $L$ by adjoining the class of the dual vector
@@
12. Proof of Theorem 1.6
If $S$ has $\chi(\mathcal{O}_{S})>0$, then the statement of Theorem 1.6 that $\mathrm{Aut}_{\mathbb{Z}}(S)$ is abelian follows from the results of this paper:
  • if $p_{g}>0$, then the statement follows from Theorem 1.1 (i);
  • if $p_{g}=0$, but the fibration is not isotrivial, then the statement can be derived from Corollary 6.3 since this implies that (8.1) also holds in this case, as in the proof of Theorem 1.1 (i);
  • if the fibration is isotrivial, then $|\mathrm{Aut}_{\mathbb{Z}}(S)|\leq 4$ by Theorem 3.6 which a posteriori implies that $\mathrm{Aut}_{\mathbb{Z}}(S)$ is abelian.
If $S$ has $\chi(\mathcal{O}_{S})=0$, we write $S$ as a free quotient of a product, $S=(C\times E)/G$. Then [CFGLS24, Thm 1.2] states explicitly that $\mathrm{Aut}_{\mathbb{Z}}(S)$ is abelian except for a special case where $G\cong\mathbb{Z}/2m\mathbb{Z}$ for odd $m$ acts by translations on $E$ and $\mathrm{Aut}_{\mathbb{Z}}(S)/\mathrm{Aut}^{0}(S)\cong\mathbb{Z}/2\mathbb{Z}=\langle\imath\rangle$. But here, the action of $\mathrm{Aut}^{0}(S)=E$ is induced by translations on the second factor of $C\times E$ while $\imath$ can only be induced by an involution on the first factor, possibly composed with translation by a 2-torsion point on $E$. In either case, these automorphisms commute on $C\times E$ and thus also on $S$. ∎
Remark 12.1.
For $\mathrm{Aut}_{\mathbb{Q}}(S)$, we only exhibited abelian examples in this paper (with $\chi>0$), but there are also non-abelian examples since one can show that any finite subgroup $H$ of $\mathrm{Aut}(\mathbb{P}^{1})$ can be realized as $\mathrm{Aut}_{\mathbb{Q}}(S)$ in case $\chi=0$.
It suffices to define $C$ as the fibre product of $B:=\mathbb{P}^{1}\rightarrow B^{\prime}:=\mathbb{P}^{1}/H$ and of a general ramified double cover $C^{\prime}\rightarrow B^{\prime}=(\mathbb{P}^{1}/H)$; letting $\iota$ be the involution of $C$ induced by the identity of $B$ and the covering involution of $C^{\prime}$, $H\times\langle\iota\rangle$ acts on $C$. We let $E=\mathbb{C}/\mathbb{Z}\oplus\tau\mathbb{Z}$ be an elliptic curve, and $S:=(C\times E)/(\mathbb{Z}/2)$, where $\mathbb{Z}/2$ acts via $(x,z)\mapsto(\iota(x),z+1/2)$. Then $H\times\{\mathrm{id}_{E}\}$ induces an inclusion $H<\mathrm{Aut}_{\mathbb{Q}}(S)$.